OpenStudy (anonymous):

an electrician charges $35 per hour plus a fixed amount for residential work. for a job that took 6 hours , he billed $270. write an equation in the form y=mx+b that represent the cost c of hiring this electrician for a job that takes x hours

OpenStudy (anonymous):

$270=35(6)+60, y=mx+b